The demand and supply functions of a goods are given by:
p=-4QD+120
p=1/2QS+28
where P, QD, and QS denote the price, quantity demanded and quantity supplied respectively. Find the equilibrium price and quantity.
1
Expert's answer
2020-12-15T02:22:57-0500
"QD=30-\dfrac{1}{4}p"

"QS=2p-56"

In equilibrium Qd = Qs


"30-\dfrac{1}{4}P=2p-56"

"\dfrac{9}{4}P=86"

"P=38.22"

"Qd=Qs=22"